### v2

# Exposure Completing for Temporally Consistent Neural High Dynamic Range Video Rendering

Abstract.
High dynamic range (HDR) video rendering from low dynamic range (LDR) videos where frames are of alternate exposures encounters significant challenges, due to the exposure change and absence at each time stamp. The exposure change and absence make existing methods generate flickering HDR results. In this paper, we propose a novel paradigm to render HDR frames via completing the absent exposure information, hence the exposure information is complete and consistent. Our approach involves interpolating neighbor LDR frames in the time dimension to reconstruct LDR frames for the absent exposures. Combining the interpolated and given LDR frames, the complete set of exposure information is available at each time stamp. This benefits the fusing process for HDR results, reducing noise and ghosting artifacts therefore improving temporal consistency. Extensive experimental evaluations on standard benchmarks demonstrate that our method achieves state-of-the-art performance, highlighting the importance of absent exposure completing in HDR video rendering. The code is available at https://github.com/cuijiahao666/NECHDR.
HDR video, Temporal consistency, Exposure completing
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680935isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Image and video acquisitionccs: Computing methodologies Computational photography
(a) Indoor scene with saturation and motions.
(b) Outdoor scene with noise and motions.
Figure 1. Qualitative comparison between the proposed method and HDRFlow (Xu et al., 2024) on two different scenes. The input low dynamic range (LDR) videos consist of frames with alternate exposures, meaning that the exposure changes and part of exposure information is missing at every time stamp ($1$-st row). Prior arts, e.g., the HDRFlow, struggle to achieve temporally consistent high dynamic range (HDR) results due to the exposure change and the motion that causes information loss ($2$-nd row). In contrast, the proposed method achieves temporally consistent HDR reconstruction results ($4$-th row) by completing the absent exposure frames ($3$-rd row) at each time stamp. “EV”: the exposure value for each input LDR frame. “EC”: our exposure completing results for the absent exposure corresponding to the input LDR frames.
[…]
1. Introduction
Compared with the high dynamic range (HDR) of natural scenes, the standard digital camera can only capture low dynamic range (LDR) information. To achieve the goal of “what you see is what you get” in computational photography (Gu et al., 2019; Pan et al., 2021, 2023), delicate optical systems (Nayar and Mitsunaga, 2000; Tocci et al., 2011; Kronander et al., 2013) are designed to simultaneously capture LDR images with different exposures which cover the whole dynamic range to reconstruct HDR results. However, these systems are expensive and un-portable, making HDR contents not easily accessible for general users (Kang et al., 2003). To address this limitation, computational methods (Kang et al., 2003; Kalantari et al., 2013; Mangiat and Gibson, 2011) are designed to render HDR videos from LDR videos whose frames are exposed alternately with different exposures (as shown in $1$-st row in Fig. 1(b)).
In this setting, the incomplete exposure information varies along the time dimension. Existing neural HDR rendering methods (Kalantari and Ramamoorthi, 2019; Chen et al., 2021; Chung and Cho, 2023; Xu et al., 2024) align neighbor LDR frames according to the LDR frame of current time stamp (called reference frame in this paper) and fuse the aligned neighbor LDR frames into the reference frame to supplement the missing exposure information. The exposure of reference frame changes at every time stamp, which means that the reference frames of different exposures may have different defects: saturation for highly exposed LDR frames, and noise for lowly exposed ones. Due to that the methods mentioned above heavily depend on reference frames, HDR results from these methods may inherit defects of the reference frames, suffering from the artifacts, i.e., the ghosting ($2$-nd row in Fig. 1(b)(a)) and noise ($2$-nd row in Fig. 1(b)(b)), thereby causing temporal inconsistency.
To tackle this problem, we revisit early traditional methods (Kang et al., 2003; Kalantari et al., 2013) for HDR rendering and are motivated to complete the missing exposure information for each time stamp (Kang et al., 2003; Kalantari et al., 2013). This idea follows the philosophy of optical HDR systems (Tocci et al., 2011) that are designed to obtain different exposed LDR frames at exact the same time. In this way, defects of different exposed frames can be mutually compensated. In this paper, we propose the Neural Exposure Completing HDR (NECHDR) framework to reconstruct the LDR frames with missing exposure information. In this way, full exposure information at each time stamp can be covered by the reference and completed LDR frames, therefore the exposure information is complete and consistent along the time dimension, which benefits the temporal consistency of the rendered HDR videos.
In the proposed NECHDR framework, pyramid features of input LDR frames are extracted by the feature encoder, then fed into the exposure completing decoder and the HDR rendering decoder. The exposure completing decoder interpolates the features of neighbor LDR frames at every level of the feature pyramid. The interpolated LDR features are combined with the features from input LDR images as the inputs to HDR rendering decoder. The HDR rendering decoder estimates coarse HDR results and the optical flows at every level. The flows can facilitate feature interpolating in exposure completing decoder. At the end of NECHDR, a simple blending network is used to integrate interpolated LDR frames, input frames, and coarse HDR frames, which can achieve high-quality and temporally consistent HDR reconstruction results.
Extensive experimental results on multiple public benchmarks demonstrate the superiority of the proposed NECHDR: by completing the missing exposure information ($3$-rd row in Fig. 1(b)), our method mitigates ghosting resulting from large motions for the time stamps with highly exposed reference frames ($4$-th row in Fig. 1(b)(a)), and reduces the noise level for the lowly exposed reference frames ($4$-th row in Fig. 1(b)(b)), hence achieves better temporal consistency. Our work sheds light again on the exposure completing for HDR video rendering. The contributions can be summarized as follows:
$\bullet$ Our work firstly implements the idea of exposure completing for neural HDR rendering.
$\bullet$ Our work proposes a novel HDR video rendering framework, a.k.a., the NECHDR, which completes the missing exposure information by interpolating LDR frames.
$\bullet$ Our NECHDR achieves new state-of-the-art performance on current benchmarks.
Figure 2. Pipline of the proposed NECHDR network. Our network mainly consists of three processes: feature encoding for LDRs, exposure completing for LDR frames with missing exposures, and HDR rendering. We extract pyramid features from the input LDR frames using a parameter shared feature encoder. Then, the optical flows predicted by the lower-level HDR rendering decoder is used to warp neighbor frames to the reference frame. Subsequently, the exposure completing decoder performs exposure completing based on the warped neighbor frame features. The input features and the completed feature are fed into the HDR rendering decoder to render the coarse HDR frame. Finally, the original frames, the completed frame, and the coarse HDR frame are fused together through a simple blending network to produce a high-quality HDR result.
[…]
pipline.
2. Related Work
2.1. HDR Image Rendering
HDR imaging technology aims to extend the dynamic range of images by merging LDR images with different exposures. Both the HDR image and video reconstructing tasks encounter the misalignment between LDR images. Pioneer studies for HDR image rendering employ image alignment techniques to address this issue. Early works (Oh et al., 2014; Khan et al., 2006; Heo et al., 2010; Zhang and Cham, 2011; Lee et al., 2014; Grosch et al., 2006) often align LDR images globally, then detect and discard unaligned regions.
Directly discarding these regions leads to significant information loss, posing challenges for HDR image rendering. To address these issues, optical-flow-based (Zimmer et al., 2011; Jacobs et al., 2008; Bogoni, 2000) and patch-based methods (Sen et al., 2012; Hu et al., 2013; Ma et al., 2017) are proposed. However, large motion still makes these methods fail to produce artifact-free HDR results. With the rapid development of deep learning, many studies for HDR image rendering switch their focus to neural networks (Niu et al., 2021; Catley-Chandar et al., 2022; Liu et al., 2023; Chung and Cho, 2022; Yan et al., 2019b; Kalantari et al., 2017; Wu et al., 2018; Yan et al., 2019a; Liu et al., 2022; Yan et al., 2023), achieving promising results. Kalantari et al(Kalantari et al., 2017) propose using neural networks to align input LDR images with predicted optical flow. Wu et al(Wu et al., 2018) directly map LDR images to HDR images, thereby avoiding alignment errors. Yan et al(Yan et al., 2019a, 2020) use spatial attention to implicitly align LDR images, achieving further improvement. Liu (Liu et al., 2022) et al.introduce a ghost-free imaging model based on swin-transformer (Liu et al., 2021). Yan et al(Yan et al., 2023) utilize patch-based and pixel-based fusion to search for information to complement the reference frame from the other frames. However, these HDR image rendering methods assume a fixed exposure for reference frames, typically medium exposure, therefore is not suitable for HDR video rendering where reference frames have alternating exposures.
2.2. HDR Video Rendering
HDR video can be obtained by delicate optical systems, such as scanline exposure/ISO (Choi et al., 2017; Heide et al., 2014), internal/external beam splitter (Kronander et al., 2014; McGuire et al., 2007; Tocci et al., 2011), modulo camera (Zhao et al., 2015) and neuromorphic camera (Han et al., 2020). However, these expensive systems are hardly accessible to general users. Therefore, rendering HDR videos from LDR videos is investigated. Kang et al(Kang et al., 2003) utilize both global and local alignment for reference frames with neighbor frames. Kalantari et al(Kalantari et al., 2013) propose a patch-based optimization algorithm to reconstruct HDR video by synthesizing missing exposures for each frame. These traditional methods offer a fundamental idea for completing missing exposures, but they require iterative optimization and are prone to producing artifacts. Recently, Kalantari et al(Kalantari and Ramamoorthi, 2019) introduce an end-to-end neural network consisting of an optical flow prediction model and a fusion network. Chen et al(Chen et al., 2021) propose a coarse-to-fine alignment framework, which utilizes optical flow for coarse alignment and then employs deformable convolution (Dai et al., 2017) for fine alignment. Chung et al(Chung and Cho, 2023) employ spatial attention instead of optical flow to achieve alignment from adjacent frames to the reference frame. Xu et al(Xu et al., 2024) propose the HDR-domain loss, which uses optical flow supervised by HDR frames for LDR alignment, yielding favorable results. However, when complex situations such as large motion, saturation, and others occur simultaneously, these methods struggle to achieve precise alignment, resulting in artifacts in the final outputs. Besides, as the exposures of reference frames alternate, the reconstructed HDR frames tend to inherit the defects of reference frames, leading to flicker in the output videos. The fundamental issue lies in the lack of exposure information at every time stamp. Different from previous methods, we propose a renaissance work which focuses on completing missing exposures for the reference frame in neural HDR video rendering. Through compensating for the defects of the alternatively exposed reference frames by utilizing the completed results, our approach avoids artifacts of ghost and noise and obtains a temporally consistent HDR video result.
3. Proposed Methods
3.1. Overview
Each frame $l_{i}$ in the LDR video $L=\{l_{i}|i=1,...,n\}$ is expected to have multiple exposures $\mathbf{E}=\{\epsilon_{i}|i=1,...,z\}$ to cover the whole dynamic range of HDR video $H=\{h_{i}|i=1,...n\}$, where $n$ is the length of video, and $z$ is the number of exposures. However, in our task, only one exposure can be obtained in one certain time stamp, which creates an exposure sequence $E=\{e_{i}|i=1,...n\}$ for the LDR video, where $e_{i}=\epsilon_{(i\oslash z)+1}$. $a\oslash b$ indicates the remainder of $a$ divided by $b$.
Following previous work (Xu et al., 2024), we try to render a high-quality HDR video with LDR video under two settings of different alternate exposures: two alternate exposures ($z=2$) with three frames as input, and three alternate exposures ($z=3$) with five frames. We use the two-exposure setting as an example to demonstrate how to achieve HDR rendering. In Sec. 3.5, we introduce how to extend our approach to the three-exposure setting. The frame $l_{t}$ corresponding to the current time stamp $t$, is treated as the reference frame. $l_{t}$ is combined with its neighbors as input $\{l_{t-1},l_{t},l_{t+1}\}$. To render HDR frames, we explicitly complete frames $\widehat{l}_{t}$ with absent exposure $\epsilon_{i}$, where $i\neq(t\oslash z)+1$. Therefore, with complete exposure information, we can render a high-quality HDR video.
Our method consists of two processes: exposure completing for the frame with the absent exposure at the time stamp $t$ and coarse-to-fine HDR rendering. These two processes are coupled together through an encoder-decoder architecture, as illustrated in the Fig. 2. The feature encoder $\mathcal{E}$ takes LDR frames $\{l_{t-1},l_{t},l_{t+1}\}$ as input and generates pyramid features for each frame. As the spatial dimensions decrease and the feature channels increase, the encoder outputs feature sets $\{\Phi_{t-1},\Phi_{t},\Phi_{t+1}\}$ for input frames, each of which is a set of $k$-level pyramid features $\Phi=\{\phi^{k}|k=1,2,3,4\}$. Considering the distinct characteristics of the two processes, we devise two dedicated decoders. Initially, an exposure completing decoder $\mathcal{D}_{I}$ is employed to interpolate frame $\widehat{l}_{t}$ and its corresponding features $\widehat{\Phi}_{t}^{k}=\{\widehat{\phi}_{t}^{k}|k=1,2,3\}$. Subsequently, a coarse-to-fine HDR rendering decoder $\mathcal{D}_{R}$ takes $\{\Phi_{t-1},\Phi_{t},\Phi_{t+1}\}$ and $\widehat{\Phi}_{t}$ as input, and outputs optical flows $F_{t-1\rightarrow t}=\{f_{t-1\rightarrow t}^{k}|k=0,1,2,3\}$, $F_{t+1\rightarrow t}=\{f_{t+1\rightarrow t}^{k}|k=0,1,2,3\}$, coarse HDR features $\widehat{\Psi}_{t}=\{\widehat{\psi}_{t}^{k}|k=1,2,3\}$ and coarse HDR frame $\widehat{h}_{t}^{c}$. The optical flows here are utilized to warp $\{\Phi_{t-1},\Phi_{t+1}\}$ or $\{l_{t-1},l_{t+1}\}$ to time stamp $t$, yielding warped neighbor features $\widetilde{\Phi}_{t-1}=\{\widetilde{\phi}_{t-1}^{k}|k=1,2,3\}$, $\widetilde{\Phi}_{t+1}=\{\widetilde{\phi}_{t+1}^{k}|k=1,2,3\}$ or frames $\{\widetilde{l}_{t-1},\widetilde{l}_{t+1}\}$. The warped neighbor features $\{\widetilde{\Phi}_{t-1},\widetilde{\Phi}_{t+1}\}$ and frames $\{\widetilde{l}_{t-1},\widetilde{l}_{t+1}\}$ are fed into $\mathcal{D}_{I}$. Subsequently, the exposure completing features $\widehat{\Phi}_{t}$ from $\mathcal{D}_{I}$ are fed into the HDR rendering decoder $\mathcal{D}_{R}$, assisting it in better restoring lost details due to saturation and noise when decoding coarse HDR features $\widehat{\Psi}_{t}$ and coarse HDR frame $\widehat{h}_{t}^{c}$. This design coupling two decoders enables mutual enhancement of the two processes, thus obtaining more accurate exposure completing frame $\widehat{l}_{t}$ and better coarse HDR result $\widehat{h}_{t}^{c}$. To obtain the final HDR rendering result, we map $\{l_{t-1},l_{t},l_{t+1}\}$ and $\widehat{l}_{t}$ to the linear HDR domain. The function defining the mapping of LDR frames to the linear HDR domain is as follows:
(1) $x_{t}=l_{t}^{\gamma}/e_{t},$
where $\gamma$ is a hyperparameter and $e_{t}$ is the exposure time of $l_{t}$. Finally, the ultimate high-quality HDR rendering result $\widehat{h}_{t}$ is obtained by fusing the input and completed frames with coarse HDR frame in linear HDR domain using a simple blending network. The details of each process will be presented in the following sections.
3.2. Exposure Completing
Under the setting that input LDR frames are with alternating exposures, completing the absent exposure information is essential for rendering HDR frames. However, existing neural HDR rendering methods overlook this essential problem, making it difficult for these methods to accurately recover detailed information when motion occurs with saturation or noise simultaneously, potentially leading to artifacts. Furthermore, due to these methods heavily depend on reference frames, the alternate appearance of noise and saturation caused by the alternate exposures in reference frames affects the temporal consistency of the video. Therefore, to better address the aforementioned issues, we explicitly tackle the fundamental problem of HDR rendering by completing frames with absent exposure corresponding to reference frames.
Given the LDR frame input$\{l_{t-1},l_{t},l_{t+1}\}$ with alternate exposures $\{e_{t-1},e_{t},e_{t+1}\}$ , where the reference frame and neighbor frames have two different exposures ( $e_{t}=\epsilon_{1}$ and $e_{t-1}=e_{t+1}=\epsilon_{2}$ ), our task is to complete the frame $\widehat{l}_{t}$ with absent exposure $\epsilon_{2}$ corresponding to reference frame through interpolation.
Our exposure completing process takes the neighbor features $\{\Phi_{t-1},\Phi_{t+1}\}$ and the optical flows $\{F_{t-1\rightarrow t},F_{t+1\rightarrow t}\}$ as input. During this process, for the features with absent exposure information, the $k$-th level neighbor features $\{\phi_{t-1}^{k},\phi_{t+1}^{k}\}$ and optical flows $\{f_{t-1\rightarrow t}^{k},f_{t+1\rightarrow t}^{k}\}$ are processed to obtain completed feature $\widehat{\phi}_{t}^{k}$:
(2) $\widetilde{\phi}_{t-1}^{k},\widetilde{\phi}_{t+1}^{k}=\mathcal{W}(\phi_{t-1}^{k},f_{t-1\rightarrow t}^{k}),\mathcal{W}(\phi_{t+1}^{k},f_{t+1\rightarrow t}^{k}),$
(3) $\widehat{\phi}_{t}^{3}=\mathcal{D}^{3}_{I}\left(\left[\widetilde{\phi}_{t-1}^{3},\widetilde{\phi}_{t+1}^{3}\right]\right),$
(4) $\widehat{\phi}_{t}^{k}=\mathcal{D}^{k}_{I}\left(\left[\widetilde{\phi}_{t-1}^{k},\widetilde{\phi}_{t+1}^{k},\mathcal{U}_{2}(\widehat{\phi}_{t}^{k+1})\right]\right),$
where $\mathcal{W}\left(\cdot,\cdot\right)$ denotes using optical flow to warp neighbor features to the reference feature, $\mathcal{U}_{2}$ represents the bilinear upsampling operation with scale factor $2$, $\mathcal{D}_{I}^{k}\left(k=1,2\right)$ represents the middle levels of the exposure completing decoder $\mathcal{D}_{I}$, and []delimited-[][$\cdot$][ ⋅ ] indicates concatenation operation. In the final step of the exposure completing process, the completed frame $\widehat{l}_{t}$ is explicitly interpolated with neighbor frames $l_{t-1},l_{t+1}$ and optical flow $f_{t-1\rightarrow t}^{0},f_{t+1\rightarrow t}^{0}$ as input:
(5) $\widetilde{l}_{t-1},\widetilde{l}_{t+1}=\mathcal{W}(l_{t-1},f_{t-1\rightarrow t}^{0}),\mathcal{W}(l_{t+1},f_{t+1\rightarrow t}^{0}),$
(6) $\widehat{l}_{t}=\mathcal{D}^{0}_{I}\left(\left[\widetilde{l}_{t-1},\widetilde{l}_{t+1},\mathcal{U}_{2}(\widehat{\phi}_{t}^{1})\right]\right),$
where $\mathcal{D}_{I}^{0}$ represents the highest level of the exposure completing decoder $\mathcal{D}_{I}$.
As illustrated in the Fig. 1(b), with the above design, we have obtained realistic exposure completing result for the challenging areas with motion, saturation and noise, providing accurate and complete exposure information for subsequent HDR rendering.
3.3. Coarse-to-fine HDR Rendering
Building upon the results of above process, we designed a coarse-to-fine HDR rendering network based on completed features $\widehat{\Phi}_{t}$ and frame $\widehat{l}_{t}$. Taking the original features $\{\Phi_{t-1},\Phi_{t},\Phi_{t+1}\}$, completed features $\widehat{\Phi}_{t}$ and warped neighbor features $\{\widetilde{\Phi}_{t-1},\widetilde{\Phi}_{t+1}\}$ as input, the HDR rendering decoder $\mathcal{D}_{R}$ reconstructs HDR features $\widehat{\Psi}_{t}$ and optical flows $\{F_{t-1\rightarrow t},F_{t+1\rightarrow t}\}$ in a coarse-to-fine manner, and finally obtains the coarse HDR frame $\widehat{h}_{t}^{c}$:
(7) $\widehat{\psi}_{t}^{3},f_{t-1\rightarrow t}^{3},f_{t+1\rightarrow t}^{3}=\mathcal{D}_{R}^{4}\left(\left[\phi_{t-1}^{4},\phi_{t}^{4},\phi_{t+1}^{4}\right]\right),$
$\widehat{\psi}_{t}^{k-1},f_{t-1\rightarrow t}^{k-1},f_{t+1\rightarrow t}^{k-1}=$
(8) $\mathcal{D}^{k}\left(\left[\phi_{t}^{k},\widehat{\phi}_{t}^{k},\widehat{\psi}_{t}^{k},f_{t-1\rightarrow t}^{k},f_{t+1\rightarrow t}^{k},\widetilde{\phi}_{t-1}^{k},\widetilde{\phi}_{t+1}^{k}\right]\right),$
$\widehat{h}_{t}^{c},f_{t-1\rightarrow t}^{0},f_{t+1\rightarrow t}^{0}=$
(9) $\mathcal{D}^{1}\left(\left[\phi_{t}^{1},\widehat{\phi}_{t}^{1},\widehat{\psi}_{t}^{1},f_{t-1\rightarrow t}^{1},f_{t+1\rightarrow t}^{1},\widetilde{\phi}_{t-1}^{1},\widetilde{\phi}_{t+1}^{1}\right]\right),$
where $\mathcal{D}_{R}^{k}\left(k=2,3\right)$ stand for the $k$-th level of the HDR rendering decoder $\mathcal{D}_{R}$. The warped neighbor features $\{\widetilde{\phi}_{t-1}^{k},\widetilde{\phi}_{t+1}^{k}\}$ help $\mathcal{D}_{R}$ to identify regions with poor alignment and subsequently obtain the refined optical flow $\{f_{t-1\rightarrow t}^{k-1},f_{t+1\rightarrow t}^{k-1}\}$.
Based on the completed frame $\widehat{l}_{t}$ and the coarse HDR frame $\widehat{h}_{t}^{c}$, we combine them with original LDR frames to obtain the final HDR rendering result $\widehat{h}_{t}$. We adopted a blending network with U-net architecture from Xu et al(Xu et al., 2024). Thus, original and completed LDR frames $\{l_{t-1},l_{t},l_{t+1},\widehat{l}_{t}\}$, along with their corresponding frames $\{x_{t-1},x_{t},x_{t+1},\widehat{x}_{t}\}$ in linear HDR domain, and coarse HDR frame $\widehat{h}_{t}^{c}$ are fed into the blending network. The blending network calculates fusion weights $W=\{w_{i}|i=0,1,2,3,4\}$ for the input five HDR domain frames $\{x_{t-1},x_{t},x_{t+1},\widehat{x}_{t},\widehat{h}_{t}^{c}\}$ and obtains the final HDR rendering result $\widehat{h}_{t}$ through a weighted average based on the computed weights $W$:
(10) $\widehat{h}_{t}=\frac{w_{0}\widehat{h}_{t}^{c}+w_{1}\widehat{x}_{t}+w_{2}x_{t}+w_{3}x_{t-1}+w_{4}x_{t+1}}{\sum_{j=0}^{4}w_{j}}.$
Following Xu et al(Xu et al., 2024), we also fuse with the neighbor frames $\{x_{t-1},x_{t+1}\}$ to provide information about static regions. As depicted in Fig. 1(b), when encountering complex scenes involving both motion and saturation or noise, the fused exposure completion frames effectively provide the missing exposure information, thereby obtaining ghost-free and low-noise HDR results. In summary, based on completing the frame with missing exposure information, we finally achieve a high-quality HDR rendering process.
Table 1. Quantitative comparisons of our method with other state-of-the-art methods on the Cinematic Video dataset (Froehlich et al., 2014).The best and the second best results are highlighted in red and blue, respectively.
Methods 2-Exposure 3-Exposure
PSNR T SSIM T HDR-VDP-2 PSNR T SSIM T HDR-VDP-2
Kalantari13 (Kalantari et al., 2013) 2013-TOG 37.51 0.9016 60.16 30.36 0.8133 57.68
Kalantari19 (Kalantari and Ramamoorthi, 2019) 2019-CGF 37.06 0.9053 70.82 33.21 0.8402 62.44
Yan19  (Yan et al., 2019a) 2019-CVPR 31.65 0.8757 69.05 34.22 0.8604 $66.18$
Prabhakar (Prabhakar et al., 2021) 2021-CVPR 34.72 0.8761 68.82 34.02 0.8633 65.00
Chen (Chen et al., 2021) 2021-ICCV 35.65 0.8949 72.09 34.15 0.8847 66.81
LAN-HDR (Chung and Cho, 2023) 2023-ICCV 38.22 0.9100 69.15 35.07 0.8695 65.42
HDRFlow (Xu et al., 2024) 2024-CVPR 39.20 0.9154 70.98 36.55 0.9039 65.89
HDRFlow (Xu et al., 2024)(+Sintel) 2024-CVPR 39.30 0.9156 71.05 36.65 0.9055 66.02
Ours 2024-MM 40.59 0.9241 73.31 37.24 0.9102 68.36
Table 2. Quantitative comparisons of our method with other state-of-the-art methods on the DeepHDRVideo dataset (Chen et al., 2021). The result is the weighted average of all results from both dynamic scenes and static scenes in this dataset. The best and the second best results are highlighted in red and blue, respectively.
Methods 2-Exposure 3-Exposure
PSNR T SSIM T HDR-VDP-2 PSNR T SSIM T HDR-VDP-2
Kalantari13 (Kalantari et al., 2013) 2013-TOG 40.33 0.9409 66.11 38.45 0.9489 57.31
Kalantari19 (Kalantari and Ramamoorthi, 2019) 2019-CGF 39.91 0.9329 71.11 38.78 0.9331 65.73
Yan19  (Yan et al., 2019a) 2019-CVPR 40.54 0.9452 69.67 40.20 0.9531 68.23
Prabhakar (Prabhakar et al., 2021) 2021-CVPR 40.21 0.9414 70.27 39.48 0.9453 65.93
Chen (Chen et al., 2021) 2021-ICCV 42.48 0.9620 74.80 39.44 0.9569 67.76
LAN-HDR (Chung and Cho, 2023) 2023-ICCV 41.59 0.9472 71.34 40.48 0.9504 68.61
HDRFlow (Xu et al., 2024) 2024-CVPR 43.18 0.9510 77.11 40.45 0.9530 72.30
HDRFlow (Xu et al., 2024)(+Sintel) 2024-CVPR 43.25 0.9520 77.29 40.56 0.9535 72.42
Ours 2024-MM 43.44 0.9558 79.20 40.13 0.9550 76.98
3.4. Training Loss
We calculate the losses for the completed LDR features $\widehat{\Phi}_{t}$ and frame $\widehat{l}_{t}$, rendered HDR features $\widehat{\Psi}_{t}$ and frame $\widehat{h}_{t}$, and optical flows $\{F_{t-1\rightarrow t},F_{t+1\rightarrow t}\}$:
The losses for images. The widely adopted $\mathcal{L}_{1}$ loss is used to supervise the completed LDR frame $\widehat{l}_{t}$ and rendered HDR frame $\widehat{h}_{t}$, and is defined as follows:
(11) $\mathcal{L}^{com}_{I}=\left\|\widehat{l}_{t}-\overline{l}_{t}\right\|_{1},$
(12) $\mathcal{L}^{ren}_{I}=\left\|\mathcal{T}\big{(}\widehat{h}_{t}\big{)}-\mathcal{T}\big{(}\overline{h}_{t}\big{)}\right\|_{1},$
(13) $\mathcal{L}_{I}=\mathcal{L}^{com}_{I}+\mathcal{L}^{ren}_{I},$
where $\overline{l}_{t}$ and $\overline{h}_{t}$ are the corresponding ground truth for the completed LDR frame $\widehat{l}_{t}$ and rendered HDR frame $\widehat{h}_{t}$, respectively. The $\mathcal{T}(\cdot)$ is a widely used function to map HDR frame to the tone-mapped HDR domain, since HDR images are typically displayed after tone mapping. This simple differentiable $\mu-$law function $\mathcal{T}(\cdot)$ is defined as follows:
(14) $\mathcal{T}(h)=\frac{\log(1+\mu h)}{\log(1+h)},$
where $\mu$ is a hyperparameter.
The losses for features. Feature space geometry loss proposed in (Kong et al., 2022) is employed to ensure that the intermediate features obtained from frame interpolation and HDR reconstruction can be more effectively refined to conform to geometrically structured features. The parameter shared encoder $\mathcal{E}$ is used to obtain corresponding pyramid features $\overline{\Phi}_{t}=\{\overline{\phi}_{t}^{k}|k=1,2,3\}$ and $\overline{\Psi}_{t}=\{\overline{\psi}_{t}^{k}|k=1,2,3\}$ from the ground truth of completed LDR and rendered HDR frames. Then, the loss function for supervising the features of completed LDR and rendered HDR frames can be written as:
(15) $\mathcal{L}^{com}_{G}=\sum_{k=1}^{3}\mathcal{L}_{cen}\big{(}\widehat{\phi}_{t}^{k},\overline{\phi}_{t}^{k}\big{)},$
(16) $\mathcal{L}^{ren}_{G}=\sum_{k=1}^{3}\mathcal{L}_{cen}\big{(}\widehat{\psi}_{t}^{k},\overline{\psi}_{t}^{k}\big{)},$
(17) $\mathcal{L}_{G}=\mathcal{L}^{com}_{G}+\mathcal{L}^{ren}_{G},$
where the $\mathcal{L}_{cen}$ is census loss (Meister et al., 2018) and computed using the soft Hamming distance between census-transformed (Zabih and Woodfill, 1994) feature maps, considering 3×3 patches in a channel-by-channel manner.
The loss for optical flows. The HDR-domain alignment loss (Xu et al., 2024) is used hierarchically to supervise the learning process of optical flow, which is defined as follows:
(18) $\mathcal{L}_{F}^{t-1\rightarrow t}=\sum_{k=0}^{3}\|\mathcal{W}\big{(}\mathcal{T}\big{(}\overline{h}_{t-1}\big{)},\mathcal{U}_{2^{k}}\big{(}f_{t-1\rightarrow t}^{k}\big{)}\big{)}-\mathcal{T}\big{(}\overline{h}_{t}\big{)}\|_{1},$
(19) $\mathcal{L}_{F}^{t+1\rightarrow t}=\sum_{k=0}^{3}\|\mathcal{W}\big{(}\mathcal{T}\big{(}\overline{h}_{t+1}\big{)},\mathcal{U}_{2^{k}}\big{(}f_{t+1\rightarrow t}^{k}\big{)}\big{)}-\mathcal{T}\big{(}\overline{h}_{t}\big{)}\|_{1},$
(20) $\mathcal{L}_{F}=\big{(}1-m_{t}\big{)}\odot\big{(}\mathcal{L}_{F}^{t-1\rightarrow t}+\mathcal{L}_{F}^{t+1\rightarrow t}\big{)},$
where $\mathcal{W}\left(\cdot,\cdot\right)$ denotes using optical flow to warp neighbor frames to the reference frame, $\mathcal{U}_{s}$ represents the bilinear upsampling operation with scale factor $s$. The mask $m_{t}$ indicates well-exposed regions in reference frame. First, $l_{t}$ is converted to YCbCr color space to extract luminance $y$. Then, $m_{t}$ is defined as $\delta_{\text{low}}<y<\delta_{\text{high}}$, where $\delta_{\text{low}}$ and $\delta_{\text{high}}$ are the low and high luminance thresholds, respectively. This allows the optical flow computation to focus more on regions with poor exposure in the reference frame.
Total loss. Our total loss can be summarized as follows:
(21) $\mathcal{L}_{total}=\mathcal{L}_{I}+\alpha*\mathcal{L}_{G}+\beta*\mathcal{L}_{F}\,.$
Figure 3. Qualitative comparison in scenes with over-saturation and motion. Left: 2-Exposure scene from the Cinematic Video dataset (Froehlich et al., 2014). Right: 2-Exposure scene from the HDRVideo dataset (Kalantari et al., 2013).”EC” refers to our exposure completion results.
[…]
ghost
Figure 4. Qualitative comparison in scenes with noise and motion. Left: 2-Exposure scene from the Cinematic Video dataset (Froehlich et al., 2014). Right: 2-Exposure scene from the DeepHDRVideo dataset (Chen et al., 2021). ”EC” refers to our exposure completion results.
[…]
noise
3.5. Extension to Three Exposures
In the alternate exposure setting with three exposures, five LDR frames $\{l_{t-2},l_{t-1},l_{t},l_{t+1},l_{t+2}\}$ are used as input to the network for HDR $\widehat{h}_{t}$ rendering. The five LDR frames are with corresponding exposure sequences $\{e_{t-2},e_{t-1},e_{t},e_{t+1},e_{t+2}\}$. In this way, the reference frame has only one certain exposure $e_{t}=\epsilon_{3}$. Specifically, the first frame $l_{t-2}$ and the fourth frame $l_{t+1}$ have the same exposure $e_{t-2}=e_{t+1}=\epsilon_{1}$, while the second frame $l_{t-1}$ and the fifth frame $l_{t+2}$ also share the same exposure $e_{t-1}=e_{t+2}=\epsilon_{2}$. Therefore, the exposure completion process takes $\{l_{t-2},l_{t},l_{t+1}\}$ and $\{l_{t-1},l_{t},l_{t+2}\}$ as inputs, generates exposure completing results $\{\widehat{l}_{t}^{\epsilon_{1}},\widehat{l}_{t}^{\epsilon_{2}}\}$ through the flow-guided completing process, and obtains coarse HDR results $\widehat{h}_{t}^{c}$ through the coarse-to-fine HDR rendering process. A total of fifteen images, including the exposure completing frames and multiple original LDR frames, along with their corresponding frames in linear HDR domain, and the coarse HDR result, are fed into the blending network. This blending network calculates seven weights to fuse the seven HDR domain images in a weighted average manner and obtain the final HDR rendering result.
4. Experiments
4.1. Experimental Setup
Datasets. We utilize synthetic training data generated from the Vimeo-90K dataset (Xue et al., 2019). To adapt the Vimeo90K dataset for HDR video reconstruction, we follow prior research (Kang et al., 2003) to convert the original data into LDR sequences with alternate exposures. To create the ground truth of completed LDR frames, we also generated LDR sequences with missing exposures in the same way. Our framework is tested on the Cinematic Video dataset (Froehlich et al., 2014) and DeepHDRVideo dataset (Chen et al., 2021). The Cinematic Video dataset has two synthetic videos from indoor and outdoor scenes. The DeepHDRVideo dataset (Chen et al., 2021) contains both real-world dynamic scenes and static scenes with random global motion augmentation. The HDRVideo dataset (Kalantari et al., 2013) is employed solely for qualitative evaluation, as it lacks ground truth.
Implementation details. We implement our approach using PyTorch and conduct experiments on an NVIDIA RTX3090 GPU. We employ AdamW optimizer (Kingma and Ba, 2015) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $10^{-4}$. In our experiments, we set $\gamma$ in Eq. 1 to 2.2 and $\mu$ in Eq. 14 to $5000$. Following HDRFlow (Xu et al., 2024), we set $\delta_{\text{low}}$ to $0.2$ and $\delta_{\text{high}}$ to $0.8$. The weighting hyperparameters $\alpha$ and $\beta$ for the loss function in Eq. 21 are set to $0.01$.
Evaluation metrics. PSNRT, SSIMT and HDR-VDP-2 (Mantiuk et al., 2011) are adopted as the evaluation metrics. PSNRT and SSIMT are computed on the tone-mapped images. HDR-VDP-2 is computed with the number of pixels per visual degree set to $30$, which means the angular resolution of the image.
4.2. Comparisons with State-of-the-art
Quantitative comparisons between our and other state-of-the-art methods on the Cinematic Video (Froehlich et al., 2014) and DeepHDRVideo (Chen et al., 2021) datasets are shown in Table 7 and Table 2, respectively. Compared to state-of-the-art methods, our approach consistently achieves superior or comparable performance. Especially, our approach achieves state-of-the-art performance on Cinematic Video (Froehlich et al., 2014) dataset, outperforming the second-best method by $1.29$dB and $0.59$dB in terms of PSNRT for the $2$-exposure and $3$-exposure settings, respectively. Qualitative comparisons are shown in Fig. 3 and Fig. 4. we compare our NECHDR with the previous methods: Chen (Chen et al., 2021), LAN-HDR (Chung and Cho, 2023) and HDRFlow (Xu et al., 2024). Fig. 3 illustrates the results from scenes with saturation and motion on Cinematic Video dataset (Froehlich et al., 2014) and HDRVideo dataset (Kalantari et al., 2013) under the $2$-exposure setting. Apart from our method, other methods tend to exhibit severe artifacts or detail loss when saturation and motion occur simultaneously. In such challenging scenarios, we achieve accurate and artifact-free exposure completion results for saturated regions by leveraging frame interpolation from neighboring frames. This enables us to fuse high-quality HDR results. And in Fig. 4, we also show the results encountering noise and motion from the Cinematic Video dataset (Froehlich et al., 2014) and DeepHDRVideo dataset (Chen et al., 2021). The scene on the left side of Fig. 4 is captured in low-light conditions, resulting in very low signal-to-noise ratio in the low-exposure frames. This can lead to very high noise levels in rendered HDR frames after tone mapping, which makes this scene particularly challenging. Specifically, methods based on optical flow (Chen et al., 2021; Xu et al., 2024) tend to produce noisy artifacts, while attention-based method (Chung and Cho, 2023) exhibit more pronounced noise. Our method explicitly reconstructs the absent high-exposure frame with low noise at the current time stamp, achieving the best noise suppression, even with some regions having noise intensity lower than ground truth. The visualization results above explain why we achieve a significant performance improvement compared with other state-of-the-art methods.
Table 3. Ablation study of NECHDR on the DeepHDRVideo (Chen et al., 2021) and CinematicVideo dataset (Froehlich et al., 2014). ”EC” refers to exposure completing. ”HR” refers to HDR rendering. ”Uncoupled” means this two processes work independently.
Model DeepHDRVideo Cinematic Video
PSNRT SSIMT PSNRT SSIMT
EC Baseline 41.09 0.9479 39.13 0.9174
HR Baseline 42.64 0.9524 39.98 0.9193
Uncoupled EC, HR 43.17 0.9535 40.27 0.9202
Coupled EC, HR (Ours) 43.44 0.9558 40.59 0.9241
Figure 5. Qualitative comparison of the models corresponding to the ablation study on a dynamic scene of DeepHDRVideo (Chen et al., 2021) dataset.
4.3. Analysis
Ablation Study. We conduct ablation experiments under the $2$-exposure setting on Cinematic Video dataset (Froehlich et al., 2014) and DeepHDRVideo dataset (Chen et al., 2021), and the quantitative and qualitative results are shown in Table 3 and Fig. 5, respectively. We devised a baseline that employs IFRNet (Kong et al., 2022) to utilize the neighbor frames to complete the middle time frame that with missing exposure information and achieve HDR results by simply fusing the completed LDR frame with the original LDR frames. The performance of this baseline is shown as “EC Baseline” in the first row in Table 3. Another baseline directly renders HDR results based on IFRNet, which is the “HR Baseline” in second row in Table 3. However, relying solely on either exposure completing or HDR rendering results is with limited performance. By adding the exposure completing decoder, we get a network (third row in Table 3) that contains both decoupled exposure completing and HDR rendering processes. Finally, through feeding the completed features and frames from exposure completing decoder into the process of HDR rendering, we achieve our NECHDR framework in the forth row in Table 3. Based on the qualitative and quantitative results, we can see: (a) coupled exposure completing and HDR rendering processes benefit the quality of final HDR results; (b) exposure completing decoder is necessary; (d) the completed results with missing exposure information help the rendering process of HDR reconstruction.
Figure 6. Visual comparisons of temporal consistency.
[…]
temporal consistency about noise
Temporal consistency. We show the visual comparisons of temporal consistency in Fig. 6. In Fig. 6, we record a two-pixel-height line traversing all frames of a scene in the Cinematic Video dataset (Froehlich et al., 2014) over time and lay them out sequentially to form temporal profiles. Base on the illustration of temporal profiles, we can observe that the horizontal stripes exist in the temporal profiles of other methods. The horizontal stripes comes from the differences between adjacent frames, which represents the temporal inconsistency. In contrast, the horizontal stripes can hardly be observed in our temporal profiles, which means that our proposed method achieves better temporal consistency.
5. Conclusion
In this paper, we implement the idea of exposure completing for neural HDR video rendering and propose the Neural Exposure Completing HDR (NECHDR) framework. The NECHDR leverages interpolation of neighbor LDR frames to complete missing exposures, providing a complete set of exposure information for each time stamp. This process of exposure completing creates a novel neural HDR video rendering pipeline, which can generate results of less noise and ghosting artifacts, thereby enhancing temporal consistency. Experimental results on multiple public benchmarks demonstrate the superiority of our NECHDR, which may shift the focus of researchers in this area to the exposure completing.
Appendix A More quantitative results
In Table 4 and Table 5, we provide the results of the current and proposed methods on the more recent metrics under the two-exposure setting, which are all tested on the CinematicVideo dataset (Froehlich et al., 2014). As shown in the Table 4, our method has achieved the state-of-the-art performance on the more recent image quality metrics. We also conduct comparisons on the video quality metric ForVideoVDP in Table 5, which considers information in temporal domain. Our method achieved state-of-the-art performance on this metric also.
Table 4. Additional results on more recent metrics
[…]
Chen 9.22 34.24 0.9190 34.45 0.9188
LAN-HDR 9.05 32.30 0.9017 33.11 0.9041
HDRFlow 9.28 33.24 0.9146 33.47 0.9168
Ours 9.30 34.66 0.9292 34.84 0.9290
Table 5. The performance on video metrics: ForVideoVDP
Methods Chen LAN-HDR HDRFlow Ours
ForVideoVDP 8.9950 8.3404 8.8692 9.0680
Appendix B The efficiency of the proposed methods
In Table 6, we provide a comparison of the efficiency between our proposed method and the current state-of-the-art methods on the CinematicVideo dataset (Froehlich et al., 2014) under the two-exposure setting. In both settings, we offer lightweight versions of the NECHDR network by reducing the number of feature channels. In the lightweight versions, feature channels from the first to the fourth levels in feature encoder $\mathcal{E}$ is set to 24, 36, 54, 72, respectively. The channels of both the exposure completing decoder $\mathcal{D}_{I}$ and HDR rendering decoder $\mathcal{D}_{R}$ are adjusted accordingly. Furthermore, the number of feature channels in the third and fifth convolution layers of HDR rendering decoder $\mathcal{D}_{R}$ will be adjusted to 24. As shown in Table 7, reducing the number of feature channels significantly decreases the parameter and inference time of our model, with no noticeable change in performance.
Table 6. The efficiency of current and proposed methods
Methods Flops(G) Para.(M) Time(ms) PSNRT SSIMT
Chen - 6.1 522 35.65 0.8949
LAN-HDR 3046 7.3 415 38.22 0.9100
HDRFlow 367 3.3 55 39.30 0.9156
Ours 1376 8.8 178 40.59 0.9241
Table 7. Quantitative comparisons with the lightweight version methods “NECHDR-lw”.
two-Exposures three-Exposures
NECHDR NECHDR-lw NECHDR NECHDR-lw
Para. (M) 8.8 4.2 ($\downarrow 52\%$) 9.9 4.8 ($\downarrow 52\%$)
Time (s) 0.178 0.138($\downarrow 23\%$) 0.163 0.130 ($\downarrow 20\%$)
PSNRT 40.59 40.33 ($\downarrow 0.7\%$) 37.24 37.11 ($\downarrow 0.3\%$)
SSIMT 0.9241 0.9217 ($\downarrow 0.3\%$) 0.9102 0.9053 ($\downarrow 0.5\%$)
Appendix C The ablation study of loss functions.
The ablation study of the loss functions is provided in Table 8. A significant performance improvement occurs after adding the exposure completion loss $\mathcal{L}^{com}_{I}$, which validates our core contribution: the importance of exposure completion for HDR video rendering. The losses for features are introduced in  (Kong et al., 2022) to provide useful low-level structure information. Here, the losses can enhance the details of HDR rendering, thereby improving performance. Finally, combined with optical flow loss $\mathcal{L}_{F}$, our method achieves the final performance.
Table 8. The ablation of losses in our method
$\mathcal{L}^{ren}_{I}$ $\mathcal{L}^{com}_{I}$ $\mathcal{L}^{ren}_{G}$ $\mathcal{L}^{com}_{G}$ $\mathcal{L}_{F}$ PSNRT SSIMT
39.83 0.9213
40.46 0.9225
40.50 0.9231
40.52 0.9932
40.59 0.9941
Figure 9. Details of the feature encoder $\mathcal{E}$
Figure 10. Details of the exposure completing decoder $\mathcal{D}_{I}$
Figure 11. Details of the HDR rendering decoder $\mathcal{D}_{R}^{1}$. Left: two-exposure; right: three-exposure.
Figure 12. Details of the HDR rendering decoder $\mathcal{D}_{R}^{2}$.
Figure 13. Details of the HDR rendering decoder $\mathcal{D}_{R}^{3}$.
Figure 14. Details of the HDR rendering decoder $\mathcal{D}_{R}^{4}$.
Appendix D More Qualitative comparisons
Under the three-exposure setting, we provide more qualitative comparison results in Fig. 7 and Fig. 8 between our method and the current state-of-the-art approaches: Chen (Chen et al., 2021), LAN-HDR (Chung and Cho, 2023), HDRFlow (Xu et al., 2024). As shown in Fig. 7, we obtain the completed frames “EC1” and “EC2”. The “EC1” provides high-exposure information with high signal-to-noise ratio to assist in removing noise from the dark background of the reference frame, while “EC2” offers low-exposure foreground object details to restore details missing due to overexposure in the reference frame. Therefore, our proposed method achieves the HDR rendering result closest to ground truth. We also provide two scenarios in Fig. 8, where motions occur with saturation and noise respectively. In the left image of Fig. 8, although “EC1” is not well completed due to the large noise and motion of the first and fourth input frames, the well-completed “EC2” provides information about the saturated regions for HDR rendering, resulting in ghost-free HDR result. In the right image of Fig. 8, the completed frame “EC2” provides detailed information with low noise, thus achieving the lowest noise level in these methods. The above visualization results elucidate the performance enhancement of our method under the three-exposure setting, further proving the robustness of our approach across different exposure settings.
Appendix E Network Architecture
We provide details of the NECHDR network we proposed, mainly including the following sub-networks: feature encoder $\mathcal{E}$, exposure completing decoder $\mathcal{D}_{I}$, $k$-levels HDR rendering decoders $\{\mathcal{D}_{R}^{k}|k=1,2,3,4\}$ and blending network. Both The feature encoder $\mathcal{E}$ and exposure-completing decoder $\mathcal{D}_{I}$ are weight-shared in the two-exposure setting or the three-exposure setting. We take input LDR frames with a size of $256*256$ as example and illustrate in figures for above sub-networks. In Fig. 91011121314, the parameters for “Conv” and “Deconv”, listed from left to right, are input channels, output channels, kernel size, stride, and padding. In our network, each “Conv” is followed by a PReLU , whereas there is no activation subsequent to each ’Deconv’ layer.
In the two-exposure setting, the parameter shared feature encoder $\mathcal{E}$ in Fig. 9 outputs pyramid features for the three input LDR frames $\{l_{t-1},l_{t},l_{t+1}\}$, while for the five input LDR frames $\{l_{t-2},l_{t-1},l_{t},l_{t+1},l_{t+2}\}$ in the three-exposure setting. As for the exposure completing decoder $\mathcal{D}_{I}$ in the Fig. 10, we adopt the same structure and parameters in both settings. Particularly, in the three-exposure setting, the parameter shared decoder $\mathcal{D}_{I}$ is utilized twice, producing two interpolated frames $\{\widehat{l}_{t}^{\epsilon_{1}},\widehat{l}_{t}^{\epsilon_{2}}\}$ and theirs corresponding features $\widehat{\Phi}_{t,1}^{k}=\{\widehat{\phi}_{t,1}^{k}|k=1,2,3\},\widehat{\Phi}_{t,2}^{k}=\{\widehat{\phi}_{t,2}^{k}|k=1,2,3\}$. The details of HDR rendering decoder $\mathcal{D}_{I}^{k}$ in $k$ levels, where $k=1,2,3,4$, are shown in Fig. 11121314, respectively. In each of the Fig. 11121314, the left illustrates the $k$-th level $\mathcal{D}_{I}^{k}$ under two-exposure setting, while the right is under three-exposure setting. The details regarding the blending network can be found in (Xu et al., 2024).